A small probe is made of stainless steel and has a length of 164.1 mm when being used to monitor the temperature in a furnace.
When the probe is removed for maintenance and allowed to cool to 20.0°C the length of the probe is measured as 162.2 mm.
What is the temperature of the furnace? (to 3 s.f and in °C)
[αsteel = 16.0×10−6 K−1]
1
Expert's answer
2020-05-24T17:59:07-0400

By the definition of the linear thermal expansion we have:


"\dfrac{\Delta L}{L_0} = \alpha \Delta T = \alpha (T_{final} - T_{initial}) ,"

here, "\Delta L = L - L_0" is the increase in length of the stainless steel probe when it is heated in the furnace, "L_0 = 0.1622 m" is the initial length of the stainless steel probe at "T_{initial} = 20 \ ^{\circ}C", "L = 0.1641 m" is the length of the stainless steel probe at temperature "T_{final}", "\alpha = 16.0 \cdot 10^{-6} \ ^{\circ}C^{-1}" is the linear expansion coefficient for steel, "\Delta T" is the change in temperature, "T_{initial} = 20 \ ^{\circ}C" is the initial temperature of the stainless steel probe, "T_{final}" is the final temperature of the stainless steel probe (or the temperature of the furnace).

Then, from this formula we can find the temperature of the furnace:


"T_{final} = \dfrac{L - L_0}{\alpha L_0} + T_{initial},""T_{final} = \dfrac{0.1641m - 0.1622m}{16.0 \cdot 10^{-6} \ ^{\circ}C^{-1} \cdot 0.1622m} + 20 \ ^{\circ}C = 752 \ ^{\circ}C."

Answer:

"T_{final} = 752 \ ^{\circ}C."